# Limit-case admissibility for positive infinite-dimensional systems

### v1

@@ 4.2 Input admissibility for positive semigroups
Suppose that Assumption 4 holds and let $U$ be an ordered Banach space whose closed unit ball is upwards directed. Then every positive operator $B\in\mathcal{L}(U,X_{-1})$ is $\mathrm{C}$-admissible.
Proof 4.13.
Without loss of generality, we assume that $\omega_{0}(A)<0$. Let $B\in\mathcal{L}(U,X_{-1})$ be positive and let $\mathrm{T}([0,\tau],U)$ denote the space of all $U$-valued step functions on $[0,\tau]$. Let $\Phi_{\tau}$ be the input operator in (4.2) with $Z=\operatorname{Reg}$. As in the proof of Theorem 4.8, $\Phi_{\tau}$ maps $\mathrm{T}([0,\tau],U)$ into $X$.
Let $u\in\mathrm{T}([0,\tau],U)$. Then we can write $u=\sum_{k=1}^{n}u_{k}\chi_{I_{k}}$ for $u_{k}\in U$ and disjoint intervals $I_{k}\subseteq[0,\tau]$. Since $B_{U}$ is upwards directed, the norm on $X^{\prime}$ is additive on the positive cone $(X^{\prime})_{+}$ by [42, Lemma 3]. Whence, according to [37, Theorem 1.3(2)], there exist $\alpha>1$ – independent of $u$ – and $0\leq w\in U$ such that $u_{k}\leq w$ for each $k$ and $\left\lVert w\right\rVert\leq\alpha\left\lVert u\right\rVert_{\infty}$. In particular, $u(s)\leq w$ for all $s\in[0,\tau]$.
Positivity of the semigroup, the operator $B$, and the element $w\in U$ now allow us to estimate $\left\lVert\int_{0}^{\tau}T_{-1}(\tau-s)Bu(s)\;\mathrm{d}s\right\rVert$ from above by
$\left\lVert\int_{0}^{\infty}T_{-1}(\tau-s)Bw\;\mathrm{d}s\right\rVert\leq\left\lVert A_{-1}^{-1}B\right\rVert\left\lVert w\right\rVert\leq\alpha\left\lVert A_{-1}^{-1}B\right\rVert\left\lVert u\right\rVert_{\infty}.$
It follows that $B$ is $\operatorname{Reg}$-admissible due to [56, Remark 4.1.7]. The $\mathrm{C}$-admissibility is therefore true due to Proposition 4.1.
Since the closed unit ball is upwards directed if and only if the dual norm is additive on the positive cone [52, Proposition 7 and Theorem 8], Proposition 3.1 and Theorem 4.12 are dual to each other.
@@
Suppose that Assumption 4 holds and assume that $X_{+}$ is a face of $X^{\prime\prime}_{+}$. If $B$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$, then $B$ is zero-class $L^{\infty}$-admissible.
Proof 4.15.
@@
Because $X_{+}$ is a face of $X^{\prime\prime}_{+}$, also $X_{+}$ is a face of $X_{-1,+}$ by Corollary 2.15. It follows that $\Phi_{\tau}u\in X$. A rescaling argument yields that $\Phi_{\tau}u\in X$ for all $u\in L^{\infty}([0,\tau],U)$.
We give some sufficient conditions in Proposition A.3 under which a positive control operator automatically satisfies the assumptions of Theorem 4.14. In particular, zero-class $L^{\infty}$-admissibility is automatic in these situations:
@@
Suppose that Assumption 4 holds and let $U$ be an ordered Banach space and let $B\in\mathcal{L}(U,X_{-1})$ be positive.
The following assumptions are sufficient for $B$ to be zero-class $L^{\infty}$-admissible.
@@
    The space $U$ has a unit and $X_{+}$ is a face of $X^{\prime\prime}_{+}$,
@@
    The unit ball of $U$ is upwards directed and every increasing norm-bounded net in $X_{+}$ is norm convergent.
Proof 4.17.
In each case, $X_{+}$ is a face of $X^{\prime\prime}_{+}$, so by Theorem 4.14, we only need to show that $B$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$. For (a), this is a direct consequence of Proposition A.3(a).
(b) Due to Corollary 2.14(a), every increasing norm-bounded net in $X_{-1,+}$ is norm-convergent. The claim now follows by Proposition A.3(b).
Here too, the example [32, Example 2.3] by Kato mentioned in Remark 4.11 shows that the assumption that $U$ has a unit cannot be dropped in Corollary 4.16(a).
4.3 Input admissibility for general semigroups
In this section, we leave the setting of Assumption 4 and instead impose order assumptions on the input space. This allows us to give conditions under which admissibility of the control operator automatically gives zero-class $L^{\infty}$-admissibility. First, we obtain the following analogue to Theorem 3.13:
@@ 4.3 Input admissibility for general semigroups
Proof 4.19.
Let $B$ be $L^{\infty}$-admissible. Then by [55, Theorem 6.9(iii)], we know that $B^{\prime}$ is $L^{1}$-admissible. Since $U$ is an AM-space, we have that $U^{\prime}$ is an AL-space. Moreover, as $X$ is reflexive, so is $X^{\prime}$. Thus, Theorem 3.13 implies that $B^{\prime}$ is also zero-class $L^{1}$-admissible. Keeping in mind that $X$ is reflexive, the result follows by [55, Theorem 6.9(ii)].
@@
For the proof of the theorem, we need the following lemma in the spirit of Weiss duality results [55, Theorem 6.9]:
@@
Proof 4.22.
Let $\tau>0$. Since $B$ is $\mathrm{C}$-admissible, there exists $K_{B,\tau}>0$ such that
$\left\lVert\int_{0}^{\tau}T_{-1}(t)Bu(t)\;\mathrm{d}t\right\rVert\leq K_{B,\tau}\left\lVert u\right\rVert_{\infty}$
for all $u\in\mathrm{C}([0,\tau],U)$. Finite-dimensionality of the input space allows us to compute
$\left\lVert B^{\prime}T^{\prime}(\mathord{\,\cdot\,})x^{\prime}\right\rVert_{1}$ $=\sup_{u\in\mathrm{C}([0,\tau],U),\left\lVert u\right\rVert_{\infty}\leq 1}\left\lvert\int_{0}^{\tau}\left\langle B^{\prime}T^{\prime}(t)x^{\prime}\,,\,u(t)\right\rangle\right\rvert\;\mathrm{d}t$
$=\sup_{u\in\mathrm{C}([0,\tau],U),\left\lVert u\right\rVert_{\infty}\leq 1}\left\lvert\int_{0}^{\tau}\left\langle B^{\prime}(T_{-1})^{\prime}(t)x^{\prime}\,,\,u(t)\right\rangle\right\rvert\;\mathrm{d}t$
$=\sup_{u\in\mathrm{C}([0,\tau],U),\left\lVert u\right\rVert_{\infty}\leq 1}\left\lvert\left\langle x^{\prime}\,,\,\int_{0}^{\tau}T_{-1}(t)Bu(t)\;\mathrm{d}t\right\rangle\right\rvert$
$\leq K_{B,\tau}\left\lVert x^{\prime}\right\rVert$
for all $x^{\prime}\in\operatorname{dom}A^{\prime}$. This proves that $B^{\prime}$ is $L^{1}$-admissible.
Proof 4.23 (Proof of Theorem 4.20).
Let $B$ be $\mathrm{C}$-admissible. Since $X$ is reflexive and $U$ is finite-dimensional, we have that $B^{\prime}$ is $L^{1}$-admissible by Lemma 4.21. It follows from Theorem 3.13 that $B^{\prime}$ is also zero-class $L^{1}$-admissible. Once again, using the reflexivity of $X$, it follows that $B$ is zero-class $L^{\infty}$-admissible by the Weiss duality result [55, Theorem 6.9(ii)].
4.4 $L^{r}$-input admissibility of positive semigroups
We close this section by moving a bit away from the limit-cases of $L^{\infty}$- and $C$-admissibility and instead considering $L^{r}$-admissibility for $r<\infty$. If one has a sufficiently strong estimate on the resolvent one can even get $L^{1}$-admissibility of positive input operators [19, Theorems 2.1]; compare however Appendix B at the end of this paper. In the following corollary, we work with an intrinsic ultracontractivity type assumption on the semigroup instead which was shown in Theorem 2.18(b) to imply that $\operatorname{span}X_{-1,+}$ is contained in a negative Favard space.
@@ 4.4 $L^{r}$-input admissibility of positive semigroups
Let $(T(t))_{t\geq 0}$ be a positive and analytic semigroup on a reflexive ordered Banach space $X$. Let $\varphi\in X^{\prime}$ be strictly positive, assume that $\operatorname{dom}A^{\prime}\cap(X^{\prime})_{\varphi}$ is majorizing in $(X^{\prime})_{\varphi}$, that $T(t)^{\prime}X^{\prime}\subseteq(X^{\prime})_{\varphi}$ for each $t>0$ and that the ultracontractivity type estimate $\left\lVert T(t)^{\prime}\right\rVert_{X^{\prime}\to(X^{\prime})_{\varphi}}\leq\widetilde{c}t^{-\alpha}$ is satisfied for constants $\widetilde{c}\geq 0$ and $\alpha\in[0,1)$ and for all $t\in(0,1]$.
If $U$ is an ordered Banach space and $B\in\mathcal{L}(U,X_{-1})$ is positive (or, more generally, the difference of two positive operators), then $B$ is $L^{r}$-admissible for every $r>\frac{1}{1-\alpha}$.
Proof 4.25.
By Theorem 2.18(b), $B$ maps into the Favard space $F_{-\alpha}$, whence, by the closed graph theorem, $B\in\mathcal{L}(U,F_{-\alpha})$. This readily implies that $B$ is $L^{r}$-admissible for $r>\frac{1}{1-\alpha}$, see e.g. [45, Lemma 2.1 and Remark 2.3] or [35, Proposition 19].
@@
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary, let $p,q\in(1,\infty)$ be Hölder conjugates and let $A$ be the elliptic operator with Neumann boundary conditions on $L^{p}(\Omega)$ from Example 2.20. If $\alpha:=\frac{d}{2q}<1$, then it follows from assertion (a) in Example 2.20 that $\operatorname{span}L^{p}(\Omega)_{-1,+}$ is contained in the Favard space $F_{-\alpha}$.
Now let $U$ be an ordered Banach space and let $B\in\mathcal{L}(U,X_{-1})$ be positive or, more generally, the difference of two positive operators. Than $B$ is $L^{r}$-admissible for every $r\in(\frac{1}{1-\alpha},\infty]$. If $s\in[1,\infty)$ denote the Hölder conjugate of $r$, this can be reformulated by saying that $B$ is $L^{r}$-admissible if $s<\frac{2q}{d}=\frac{1}{\alpha}$.
We note that the derived range of $r$ may not be optimal, partly due to its dependence on the dimension $d$. Indeed, it is known that the operator $B$ given by Neumann boundary control is $L^{r}$-admissible for $r>4/3$, if the boundary of $\Omega$ is $C^{2}$, see e.g. [34] or [48, Example 2.14] and even $r\geq 4/3$ when the boundary is $C^{\infty}$ [27, Proposition 2.4] and [45, Remark 2.8].
5 Applications to perturbation results
The consequences of admissibility for perturbation results are well-known. Indeed, if $(T(t))_{t\geq 0}$ is a $C_{0}$-semigroup on a Banach space and $B\in\mathcal{L}(X,X_{-1})$ is zero-class $\mathrm{C}$-admissible, then the restriction of $A_{-1}+B$ to $X$ generates a $C_{0}$-semigroup on $X$ [17, Corollary III.3.3]. Likewise, zero-class $L^{1}$-admissibility of the observation operator also yields a perturbation result [17, Corollary III.3.16]. Using results obtained in the prequel, we are thus able to obtain perturbation results for positive $C_{0}$-semigroups.
@@ 5 Applications to perturbation results
Proof 5.2.
The assumptions imply, according to Theorem 4.8, that $B$ is zero-class $\mathrm{C}$-admissible. Hence, the claim follows from the perturbation result in [17, Corollary III.3.3].
As a consequence of Corollary 5.1, one can at once see that the operator $A_{-1}+B$ in [10, Example 5.1] generates a $C_{0}$-semigroup without having to check any spectral conditions and one can also allow multiplication operator with function in $L^{1}$ as perturbations:
@@
generates a positive $C_{0}$-semigroup [17, Example II.3.19(i)] that corresponds to the partial differential equation
@@
[…]
@@
The perturbation assertions of Corollary 5.4 hold once we establish that $C$ is zero-class $L^{1}$-admissible [17, Corollary III.3.16]. Let us prove this separately in the spirit of Section 3.2:
@@
Proof 5.6.
Since $X$ is reflexive, $X_{+}$ is a face of $X^{\prime\prime}_{+}$. Therefore, $C^{\prime}\in\mathcal{L}(Y^{\prime},X^{\prime}_{-1})$ is zero-class $L^{\infty}$-admissible by Theorem 4.14. Zero-class $L^{1}$-admissibility of $C$ is now a consequence of Weiss duality result [55, Theorem 6.9(a)].
Note that this theorem implies in particular the following result: every positive finite rank perturbation $C:X_{1}\to X$ (and more generally, every finite rank perturbation that can be written as the difference of two positive operators $X_{1}\to X$) of a positive $C_{0}$-semigroup on a reflexive Banach lattice again generates a $C_{0}$-semigroup. In fact, this is known to be true even if $X$ is not reflexive and goes back to Arendt and Rhandi [8, Corollary 2.4]. Note that this is not true, in general, for finite rank perturbations that do not satisfy any positivity assumption: Desch and Schappacher proved in [15, Theorem 2] that a $C_{0}$-semigroup on a Banach space $X$ is automatically analytic if every rank-$1$ perturbation $C:X_{1}\to X$ still generates a $C_{0}$-semigroup.

### v4

@@ 4.2. Input admissibility for positive semigroups
Suppose that Assumption 4.1 holds and $U$ is an ordered Banach space whose open unit ball is upwards directed. If $B\in\mathcal{L}(U,X_{-1})$ is positive, then it is $\mathrm{C}$-admissible.
Proof.
Without loss of generality, we assume that $\omega_{0}(A)<0$. Let $\Phi_{\tau}$ be the input operator defined in (4.2) with $Z=\operatorname{Reg}$. As in the proof of Theorem 4.6, $\Phi_{\tau}$ maps the space of step functions $\mathrm{T}([0,\tau],U)$ into $X$.
Let $u\in\mathrm{T}([0,\tau],U)$. Then we can write $u=\sum_{k=1}^{n}u_{k}\chi_{I_{k}}$ for $u_{k}\in U$ and disjoint intervals $I_{k}\subseteq[0,\tau]$. Since the inerior of $B_{U}$ is upwards directed, the norm on $U^{\prime}$ is additive on the positive cone $(U^{\prime})_{+}$ by [48, Lemma 3]. Whence, according to [43, Theorem 1.3(2)], there exist $\alpha>1$ – independent of $u$ – and $w_{1},w_{2}\in U$ such that $-u_{k}\leq w_{1}$ and $u_{k}\leq w_{2}$ for each $k$ and $\left\lVert w_{i}\right\rVert\leq\alpha\left\lVert u\right\rVert_{\infty}$. In particular, $-w_{1}\leq u(s)\leq w_{2}$ for all $s\in[0,\tau]$.
Due to normality of the cone, we can choose $c>0$ such that (4.3) holds. Positivity of the semigroup, the operator $B$, and the elements $w_{1},w_{2}\in U_{+}$ now allow us to estimate $\left\lVert\int_{0}^{\tau}T_{-1}(\tau-s)Bu(s)\;\mathrm{d}s\right\rVert$ from above by
$c\max_{i=1,2}\left\lVert\int_{0}^{\infty}T_{-1}(\tau-s)Bw_{i}\;\mathrm{d}s\right\rVert\leq c\left\lVert A_{-1}^{-1}B\right\rVert\max_{i=1,2}\left\lVert w_{i}\right\rVert\leq c\ \alpha\left\lVert A_{-1}^{-1}B\right\rVert\left\lVert u\right\rVert_{\infty}.$
It follows that $B$ is $\operatorname{Reg}$-admissible by [62, Remark 4.1.7]. The $\mathrm{C}$-admissibility is therefore true due to Proposition 4.2. ∎
Since the open unit ball is upwards directed if and only if the dual norm is additive on the positive cone [58, Proposition 7 and Theorem 8], Proposition 3.2 and Theorem 4.9 are dual to each other.
@@
Suppose that Assumption 4.1 holds and assume that $X_{+}$ is a face of $X^{\prime\prime}_{+}$. If $B$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$, then $B$ is zero-class $L^{\infty}$-admissible.
Proof.
@@
Because $X_{+}$ is a face of $X^{\prime\prime}_{+}$, also $X_{+}$ is a face of $X_{-1,+}$ by Corollary 2.10. It follows that $\Phi_{\tau}u\in X$. A rescaling argument yields that $\Phi_{\tau}u\in X$ for all $u\in L^{\infty}([0,\tau],U)$. ∎
Assumptions analogous to Theorem 4.10 were recently used to provide a duality result in [12, Theorem 4.4]. From Theorem 4.10 and Proposition A.2 we get that zero-class $L^{\infty}$-admissibility is automatic in the following situations.
@@
Let Assumption 4.1 hold, assume that $U$ is an ordered Banach space, and let $B\in\mathcal{L}(U,X_{-1})$ be positive. Each of the following assumptions is sufficient for $B$ to be zero-class $L^{\infty}$-admissible.
@@
    The space $U$ has a unit and $X_{+}$ is a face of $X^{\prime\prime}_{+}$.
@@
    The open unit ball of $U$ is upwards directed and $X$ is a KB-space.
Proof.
In each case, $X_{+}$ is a face of $X^{\prime\prime}_{+}$ [63, Theorem 7.1], so by Theorem 4.10, we only need to show that $B$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$. For (a), this is shown in Proposition A.2(a).
(b) Due to Corollary 2.9(a), every increasing norm-bounded net in $X_{-1,+}$ is norm-convergent. The claim now follows by Proposition A.2(b). ∎
Here again, the example [37, Example 2.3] by Kato mentioned in Remark 4.8 shows that the assumption that $U$ has a unit cannot be dropped in Corollary 4.11(a).
4.3. Input admissibility for general semigroups
In this section, we leave the setting of Assumption 4.1 and instead impose order assumptions on the input space. This allows us to give conditions under which admissibility of the control operator automatically gives zero-class $L^{\infty}$-admissibility. First, we obtain the following analogue to Theorem 3.9:
@@ 4.3. Input admissibility for general semigroups
Proof.
Since $B$ be $L^{\infty}$-admissible, $B^{\prime}$ is $L^{1}$-admissible [61, Theorem 6.9(iii)]. Since $U$ is an AM-space, we have that $U^{\prime}$ is an AL-space. Moreover, as $X$ is reflexive, so is $X^{\prime}$. Thus, Theorem 3.9 implies that $B^{\prime}$ is zero-class $L^{1}$-admissible. Keeping in mind that $X$ is reflexive, the result follows by [61, Theorem 6.9(ii)]. ∎
@@
The proof uses the following duality lemma in the spirit of Weiss [61, Theorem 6.9]:
@@
Proof.
Let $\tau>0$. Since $B$ is $\mathrm{C}$-admissible, there exists $K_{B,\tau}>0$ such that $\left\lVert\int_{0}^{\tau}T_{-1}(t)Bu(t)\;\mathrm{d}t\right\rVert\leq K_{B,\tau}\left\lVert u\right\rVert_{\infty}$ for all $u\in\mathrm{C}([0,\tau],U)$. Finite-dimensionality of the input space allows us to compute
$\left\lVert B^{\prime}T^{\prime}(\mathord{\,\cdot\,})x^{\prime}\right\rVert_{1}$ $=\sup_{u\in\mathrm{C}([0,\tau],U),\left\lVert u\right\rVert_{\infty}\leq 1}\left\lvert\int_{0}^{\tau}\left\langle B^{\prime}T^{\prime}(t)x^{\prime}\,,\,u(t)\right\rangle\;\mathrm{d}t\right\rvert$
$=\sup_{u\in\mathrm{C}([0,\tau],U),\left\lVert u\right\rVert_{\infty}\leq 1}\left\lvert\int_{0}^{\tau}\left\langle B^{\prime}(T_{-1})^{\prime}(t)x^{\prime}\,,\,u(t)\right\rangle\;\mathrm{d}t\right\rvert$
$=\sup_{u\in\mathrm{C}([0,\tau],U),\left\lVert u\right\rVert_{\infty}\leq 1}\left\lvert\left\langle x^{\prime}\,,\,\int_{0}^{\tau}T_{-1}(t)Bu(t)\;\mathrm{d}t\right\rangle\right\rvert\leq K_{B,\tau}\left\lVert x^{\prime}\right\rVert$
for all $x^{\prime}\in\operatorname{dom}A^{\prime}$. This proves that $B^{\prime}$ is $L^{1}$-admissible. ∎
Proof of Theorem 4.13.
Since $B$ be $\mathrm{C}$-admissible, $B^{\prime}$ is zero-class $L^{1}$-admissible by Lemma 4.14 and Theorem 3.9. Once again, using the reflexivity of $X$, it follows that $B$ is zero-class $L^{\infty}$-admissible by the Weiss duality result [61, Theorem 6.9(ii)]. ∎
4.4. $L^{r}$-input admissibility of positive semigroups
We close this section by moving away from the limit-cases of $L^{\infty}$- and $\mathrm{C}$-admissibility and instead considering $L^{r}$-admissibility for $r<\infty$. If one has a sufficiently strong estimate on the resolvent one can even get $L^{1}$-admissibility of positive input operators [22, Theorems 2.1]; compare however Appendix B. In the following corollary, we instead work with an intrinsic ultracontractivity type assumption on the semigroup. In Theorem 2.12(b), this same condition was shown to imply that $\operatorname{span}X_{-1,+}$ is contained in a Favard space.
@@ 4.4. $L^{r}$-input admissibility of positive semigroups
Let $(T(t))_{t\geq 0}$ be a positive and analytic $C_{0}$-semigroup on a reflexive ordered Banach space $X$ whose cone $X_{+}$ is generating. Let $\varphi\in X^{\prime}$ be strictly positive, assume that $\operatorname{dom}A^{\prime}\cap(X^{\prime})_{\varphi}$ is majorizing in $(X^{\prime})_{\varphi}$, that $T(t)^{\prime}X^{\prime}\subseteq(X^{\prime})_{\varphi}$ for each $t>0$ and that the ultracontractivity type estimate $\left\lVert T(t)^{\prime}\right\rVert_{X^{\prime}\to(X^{\prime})_{\varphi}}\leq\widetilde{c}t^{-\alpha}$ holds for some $\widetilde{c}\geq 0$ and $\alpha\in[0,1)$ and for all $t\in(0,1]$.
If $U$ is an ordered Banach space and $B\in\mathcal{L}(U,X_{-1})$ is positive (more generally, the difference of two positive operators), then $B$ is $L^{r}$-admissible for every $r>\frac{1}{1-\alpha}$.
Proof.
By Theorem 2.12(b), $B$ maps into the Favard space $F_{-\alpha}$, whence, by the closed graph theorem, $B\in\mathcal{L}(U,F_{-\alpha})$. This readily implies that $B$ is $L^{r}$-admissible for $r>\frac{1}{1-\alpha}$, see e.g. [51, Lemma 2.1 and Remark 2.3] or [41, Proposition 19]. ∎
@@
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary, let $p,q\in(1,\infty)$ be Hölder conjugates, and let $A$ be the elliptic operator with Neumann boundary conditions on $L^{p}(\Omega)$ from Example 2.13. If $\alpha:=\frac{d}{2q}<1$, then it follows from (a) in Example 2.13 that $\operatorname{span}L^{p}(\Omega)_{-1,+}$ is contained in the Favard space $F_{-\alpha}$.
Now let $U$ be an ordered Banach space and let $B\in\mathcal{L}(U,X_{-1})$ be positive or, more generally, the difference of two positive operators. Then $B$ is $L^{r}$-admissible for every $r\in(\frac{1}{1-\alpha},\infty]$ by Corollary 4.15. If $s\in[1,\infty)$ denotes the Hölder conjugate of $r$, this can be reformulated by saying that $B$ is $L^{r}$-admissible if $s<\frac{2q}{d}=\frac{1}{\alpha}$.
We note that a fixed positive operator $B$ can be $L^{r}$-admissible for a larger range of $r$. For instance, $B$ obtained from the Neumann boundary control is $L^{r}$-admissible for $r>4/3$, if the boundary of $\Omega$ is $C^{2}$, see e.g. [40] or [54, Example 2.14], and even for $r\geq 4/3$ when the boundary is $C^{\infty}$ [32, Proposition 2.4] and [51, Remark 2.8].
5. Applications to perturbation results
The consequences of admissibility for perturbation results are well-known. Indeed, if $(T(t))_{t\geq 0}$ is a $C_{0}$-semigroup on a Banach space $X$ and $B\in\mathcal{L}(X,X_{-1})$ is zero-class $\mathrm{C}$-admissible, then the restriction of $A_{-1}+B$ to $X$ generates a $C_{0}$-semigroup on $X$ [20, Corollary III.3.3]. As a consequence, perturbation results for positive $C_{0}$-semigroups on AM-spaces were proved in [13]. Likewise, zero-class $L^{1}$-admissibility of the observation operator also yields a perturbation result [20, Corollary III.3.16]. Using results obtained in the prequel, we are thus able to obtain perturbation results for positive $C_{0}$-semigroups.
@@ 5. Applications to perturbation results
Proof.
The assumptions imply that $B$ is zero-class $\mathrm{C}$-admissible (Theorem 4.6). Hence, the claim follows from the perturbation result in [20, Corollary III.3.3]. ∎
As a consequence of Corollary 5.1, one can see at once that the operator $A_{-1}+B$ in [13, Example 5.1] generates a $C_{0}$-semigroup without verifying any spectral conditions and one can also allow multiplication operator with function in $L^{1}$ as perturbations:
@@
generates a positive $C_{0}$-semigroup [20, Example II.3.19(i)] that corresponds to the partial differential equation
@@
Let $\mu$ be a finite continuous positive Borel measure on $(0,1)$ and let $m\in L^{1}(0,1)$. It can be shown using standard arguments that $\mu,m\in X_{-1}$ (see, for instance, [13, Page 344]). In particular, the rank-one operator $f\mapsto B_{1}f:=\int_{0}^{1}f(x)\;\mathrm{d}x\cdot\mu$ and the multiplication operator $f\mapsto B_{2}f:=mf$ both lie in $\mathcal{L}(X,X_{-1})$. In fact, $B_{1}$ and $B_{2}$ map the unit ball into the order-bounded subsets $[-\mu,\mu]$ and $\big{[}-\left\lvert m\right\rvert,\left\lvert m\right\rvert\big{]}$ of $X_{-1}$ respectively. It follows by Corollary 5.1 that the part of $A_{-1}+B_{1}$ and $A_{-1}+B_{2}$ in $X$ generate $C_{0}$-semigroups on $X$.
@@
The perturbation assertions of Corollary 5.3 hold once we establish that $C$ is zero-class $L^{1}$-admissible [20, Corollary III.3.16]. Let us prove this separately in the spirit of Section 3.2:
@@
Proof.
Since $X$ is reflexive, $X_{+}$ coincides with $X^{\prime\prime}_{+}$. Therefore, $C^{\prime}\in\mathcal{L}(Y^{\prime},X^{\prime}_{-1})$ is zero-class $L^{\infty}$-admissible by Theorem 4.10. Zero-class $L^{1}$-admissibility of $C$ is now a consequence of Weiss duality result [61, Theorem 6.9(a)]. ∎
Note that Corollary 5.3 implies in particular the following result: every positive finite rank perturbation $C:X_{1}\to X$ (and more generally, every finite rank perturbation that can be written as the difference of two positive operators $X_{1}\to X$) of a positive $C_{0}$-semigroup on a reflexive Banach lattice again generates a $C_{0}$-semigroup. In fact, this is known to be true even if $X$ is not reflexive and goes back to Arendt and Rhandi [8, Corollary 2.4]. Note that this is not true, in general, for finite rank perturbations that do not satisfy any positivity assumption: Desch and Schappacher proved in [18, Theorem 2] that a $C_{0}$-semigroup on a Banach space $X$ is automatically analytic if every rank-$1$ perturbation $C:X_{1}\to X$ still generates a $C_{0}$-semigroup.
Acknowledgements
The first author was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 523942381. The article was initiated during a pleasant visit of the first author to the third author at Tampere University, Finland in autumn 2023. The first author is indebted to COST Action 18232 for financial support for this visit. The research of the third author was supported by the Research Council of Finland grant number 349002.